# references.bib

### v1

@@ 3.1 Large extremal number
Let $H\in\mathrm{EX}(n,\mathcal{F})$. 1 gives that $H$ is obtained from $T_{n,r}$ by embedding trees in one part $U_{1}$, plus a bounded number of edge modifications. By (6), $e(U_{i})=0$ for any $i\neq 1$. By (2), every component of $H[U_{1}]$ is of order at most $k+1$, so by (5), every component is a tree. By rearranging the edges using (2) and (3), there is a graph $H^{\prime}$ with $e(H^{\prime})=e(H)$, such that $H^{\prime}[U_{1}]\simeq P_{k+1}\cup(\lfloor(|U_{1}|-k-1)/k\rfloor\cdot P_{k})\cup T$, where $T$ is a tree of order at most $k-1$; moreover $T$ is $\mathcal{F}$-free. (If there are at least $2$ smaller trees in the $k-1$ remaining vertices, then combining them shows that $H^{\prime}\not\in\mathrm{EX}(n,\mathcal{F})$). Since $\lfloor n/r\rfloor\equiv 0\pmod{k}$, we must have $|V_{1}|=\lceil n/r\rceil$, or else $H^{\prime}\not\in\mathrm{EX}(n,\mathcal{F})$. Now it is evident that $H[V_{1}]$ contains a $P_{k+1}$, or else $e(H)<e(H^{\prime})$. Thus by (4), $H[U_{1}]$ has no $K_{1,k-1}$. Finally, using (3) and (1), we deduce that the trees in $H[U_{1}]$ other than the $P_{k+1}$ are all $P_{k}$. Hence, $H=H$. ∎
@@ 3.2 Chromatic number 3
Let $\mathcal{F}$ consist of $K_{4}$ and all graphs of the form $T_{4}+(m\cdot K_{1})$, where $T_{4}$ is a tree on four vertices. (Here $m$ is a sufficiently large constant.) Now $\chi(F)=4$ and $\mathcal{F}$ satisfies the assumptions of 1 with $Q=\frac{1}{3}$. Let $n=2p$, where $p\equiv 1\pmod{3}$. Let $G$ be the graph obtained by embedding $U(P_{3},p-1)$ in the smaller part of $K_{p-1,p+1}$; let $H$ be the graph obtained by embedding $U(P_{3},p)$ in one part of $K_{p,p}$, and let $H^{\prime}$ be the graph obtained by embedding $U_{P_{3},p-4}\cup P_{2}\cup P_{2}$ in one part of $K_{p,p}$. After applying 1 it is easy to check that $\mathrm{EX}(n,\mathcal{F})=\{H,H^{\prime}\}$. Now $G,H,H^{\prime}$ have equitable partitions with quotient matrices
$B=\begin{pmatrix}0&2&p+1\\ 1&0&p+1\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&0\end{pmatrix},\ \ C=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&0&p\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&1&0\end{pmatrix},\ \ C^{\prime}=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&1&p\\ \frac{p-4}{3}&\frac{2(p-4)}{3}&4&0\end{pmatrix}$
respectively. We find that $\lambda(H),\lambda(H^{\prime})<p+\frac{2}{3}-\frac{1}{5p}<\lambda(G)$ by substituting this value into the characteristic polynomials. Since $G\not\in\mathrm{EX}(n,\mathcal{F})$, this implies that $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$. ∎
@@ 5 Concluding remarks
Theorem 2 gives a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which guarantees that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, when $\chi(\mathcal{F})=r+1\geq 4$. This condition is sharp for many values of $r$. As shown by 1, Theorem 2 does not hold if we allow $r=2$ and $k(\mathcal{F})=3$. We attempted to provide a similar counterexample in the case of $r=2,k(\mathcal{F})=2$ but were unable to do so because $\lambda(H)$ is in fact greater than $\lambda(G)$ by a term of $\Theta(n^{-2})$, where $G$ and $H$ are graphs defined analogously to the ones in subsection 3.2. This fact, combined with the result of [wang2023conjecture], suggested 1. A solution to 1 would be interesting as it potentially requires stronger techniques than those of this paper and those of [wang2023conjecture].
@@
Besides 1 we find the following open questions to be most interesting.

### v2

@@ 3.1 Large extremal number
Let $H^{\prime}\in\mathrm{EX}(n,\mathcal{F})$. Following subsection 2.1 through 15 shows that $H^{\prime}$ is obtained from $K_{W_{1},\ldots,W_{r}}$ by embedding trees in $W_{1}$, where $|w_{i}-w_{j}|\leq 1$ for $i,j>1$ and $w_{1}-2\leq w_{i}\leq w_{1}+1$ for $i>1$. Suppose there are $h$ parts of size $w_{1}-2$ and $r-h-1$ parts of size $w_{1}-1$, so that
$w_{1}+h(w_{1}-2)+(r-h-1)(w_{1}-1)=n\Longrightarrow w_{1}=\frac{n-1}{r}+1+\frac{h}{r}.$
Now $n\equiv 1\pmod{r}$, the right-hand side is an integer, and $10\leq h\leq r-$ 1, so we have $h=0$, and in fact $K_{W_{1},\ldots,W_{r}}=T_{n,r}$. By (1) through (5), $H^{\prime}[W_{1}]$ is a disjoint union of trees, each of which is $P_{k+1}$, $P_{k}$, $K_{1,{k-1}}$, or of order smaller than $k$. Moreover, at most one tree is $P_{k+1}$ and such a tree can only exist if there is no $K_{1,k-1}$. From these restrictions and the fact that $\lfloor n/r\rfloor\equiv 0\pmod{k}$ it is clear that $e(H^{\prime}[W_{1}])$ (hence $e(H^{\prime})$) is maximized exactly when $H^{\prime}\simeq H$. ∎
@@ 3.2 Chromatic number 3
Let $\mathcal{F}$ consist of $K_{4}$ and all graphs of the form $T_{4}+(m\cdot K_{1})$, where $T_{4}$ is a tree on four vertices. (Here $m$ is a sufficiently large constant.) Now $\chi(F)=3$ and $\mathcal{F}$ satisfies the assumptions of 1 with $Q=\frac{1}{3}$. Let $n=2p$, where $p\equiv 1\pmod{3}$. Let $G$ be the graph obtained by embedding $(p-1)/3\cdot P_{3}$ in the smaller part of $K_{p-1,p+1}$; let $H$ be the graph obtained by embedding $(p-1)/3\cdot P_{3}$ in one part of $K_{p,p}$, let $H^{\prime}$ be the graph obtained by embedding $((p-4)/3\cdot P_{3})\cup(2\cdot P_{2})$ in one part of $K_{p,p}$, and let $H^{\prime\prime}$ be the graph obtained by embedding $((p-1)/3\cdot P_{3})\cup P_{2}$ in the large part of $K_{p+1,p-1}$. Similarly to 21 one can prove $\mathrm{EX}(n,\mathcal{F})=\{H,H^{\prime},H^{\prime\prime}\}$. Now $G,H,H^{\prime},H^{\prime\prime}$ have equitable partitions with quotient matrices
$B=\begin{pmatrix}0&2&p+1\\ 1&0&p+1\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&0\end{pmatrix},$ $\ \ C=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&0&p\\ \frac{p-1}{3}&\frac{2(p-1)}{3}&1&0\end{pmatrix},$
$C^{\prime}=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&1&p\\ \frac{p-4}{3}&\frac{2(p-4)}{3}&4&0\end{pmatrix},$ $\ \ C^{\prime\prime}=\begin{pmatrix}0&2&0&p\\ 1&0&0&p\\ 0&0&1&p\\ \frac{p-4}{3}&\frac{2(p-4)}{3}&4&0\end{pmatrix},$
respectively. We find that $\lambda(H),\lambda(H^{\prime}),\lambda(H^{\prime\prime})<p+\frac{2}{3}-\frac{1}{5p}<\lambda(G)$ by substituting this value into the characteristic polynomials. Since $G\not\in\mathrm{EX}(n,\mathcal{F})$, this implies that $\mathrm{SPEX}(n,\mathcal{F})\cap\mathrm{EX}(n,\mathcal{F})=\emptyset$. ∎
@@ 5 Concluding remarks
Theorem 2 gives a condition on the value of $\mathrm{ex}(n,\mathcal{F})$ which guarantees that $\mathrm{SPEX}(n,\mathcal{F})\subseteq\mathrm{EX}(n,\mathcal{F})$, when $\chi(\mathcal{F})=r+1\geq 4$. This condition is sharp for many values of $r$. As shown by 2, Theorem 2 does not hold if we allow $r=2$ and $k(\mathcal{F})=3$. We attempted to provide a similar counterexample in the case of $r=2,k(\mathcal{F})=2$ but were unable to do so because $\lambda(H)$ is in fact greater than $\lambda(G)$ by a term of $\Theta(n^{-2})$, where $G$ and $H$ are graphs defined analogously to the ones in subsection 3.2. This fact, combined with the result of [wang2023conjecture], suggested 1. A solution to 1 would be interesting as it potentially requires stronger techniques than those of this paper and those of [wang2023conjecture].
@@
Besides 1, the open questions we are most interested in are the following.